HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025  Not logged in ELOG logo
Message ID: 64     Entry time: Wed Apr 2 05:48:31 2025
Author: Chris, Carlo, Zac 
Category: General 
Subject: Beam returned 
Beam was returned to the ESR at 05:11.

HKR said there had been some problem with one of the dipoles and in order to restore it a power supply had to be reset, which took some time.

05:41 data are back in the RSAs. Files that contain data from:
RSA01-2025.04.02.05.41.14.938.tiq_spectrogram.npz
to:
RSA01-2025.04.02.06.24.01.530.tiq_spectrogram.npz


06:24 data not visible again, only noise.

06:43 data are ack in the RSAs. Files that contain data from:
RSA01-2025.04.02.06.43.43.580.tiq_spectrogram.npz
ELOG V3.1.5-fc6679b